OpenStudy (anonymous):

A ball is dropped from the top of a 550 ft. building. The function h(t) = - 16t2 + 50 models the height of the ball, h(t) (in feet), at any given time, t (in seconds). What is the maximum height of the ball? 423 ft 550 ft

OpenStudy (danjs):

is h(t) -16t^2 + 550 instead of that 50?

OpenStudy (anonymous):

its 50

OpenStudy (raffle_snaffle):

It can't just be 50

OpenStudy (danjs):

the height function is decreasing faster and faster (acceleration) the max is when t=0, should be th etop of the building at h(0)=550ft

OpenStudy (danjs):

the number on the end of h(t) is the initial height h(0), in this case, it is the same as the building height

OpenStudy (anonymous):

I thought it was 423 ft , but i guessed and i really need to know how to do these because i have a test with like 10 problems.

OpenStudy (danjs):

is this just math class with given functions and stuff, or you in physics doing kinematics

OpenStudy (anonymous):

Its algrbra

OpenStudy (danjs):

time t, will be just positive, so the graph of h(t) is just the right half of an upside down parabola

OpenStudy (danjs):

the height starts at time t=0 h(0) = 50 vertex of the parabola, maximum value in this case

OpenStudy (danjs):

however, it says it was dropped from a height of 550, one thing is not right

OpenStudy (anonymous):

but i dont know which to put ??

OpenStudy (phi):

There may be a typo, but to make sense, the equation is probably \[ h(t) = -16t^2 + 550\] in this problem, they say the ball is dropped from the top of a building (we would say at time t=0) the ball then falls. The highest the ball will be is at the start, at 550 it then falls

OpenStudy (anonymous):

Right

OpenStudy (phi):

if we did it just using math (and using the correct equation h(t)= -16 t^2 + 0 t + 550 (I put in the missing "t" term) and match that with y= a t^2 + b t + c we match a with -16 b with 0 c with 550 and we use the formula that the vertex (highest point of the parabola) happens at t= -b/(2a) with b=0 and a= -16 we get t= -0/(2*-16) = 0 the math says the highest point is at t=0 and the highest height will be -16*0*0 +550 = 0+550= 550

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

But why add 550 ?

OpenStudy (phi):

I used the formula \[ h(t)= -16t^2 + 550\] instead of \( h(t)= -16t^2 +50\) because the new equation "works" (I know the ball is 550 ft high (on top of the building) and the second equation says the ball is at 50 feet high (at time t=0) also, I know math problems sometimes have typos. So I fixed the equation.